OpenStudy (anonymous):

Thought I'd put this in Physics too! 160 grams of ethanol at 30 degrees Celsius is mixed with 250 grams of ethanol at 77 degrees Celsius at constant atmospheric pressure in a thermally insulated vessel. What is the Delta S sys for the process? Specific heat capacity for ethanol is 2.42 J/g K. So I got up to the part of getting what t should equal 331.8085. I just don't know how to do the rest of it after Delta S sys because the teacher didn't teach us. Help?

OpenStudy (anonymous):

Aaaah Thermodynamic where Chemistry meets Physics and everybody is confused. We can only calculate change in Entropy. Also this change only depends on the initial and final state and not on how it got from one to the other no matter what the process. In your problem you have 250ml @350K and 160ml @ 303K going to 410ml @ T K which you found to be 331.8 K. This would be the same as if you took the 250ml @350K to 331.8K and took the 160ml from 303K to 331.8K. And the reason that we want to do this is that we have developed expressions for calculating the change in entropy when a given quantity of a substance is heated or cooled. That formula is\[S _{f}-S _{i}= c _{p}\ln \left( \frac{ T _{f} }{ T _{i} } \right)\]Which I assume was discussed. Calculate the change in entropy for each quantity of ethanol. One will increase (the result is positive) and one will decrease (the result is negative). ; Sum the two for your answer. Note the answer is also the difference between the absolute value of each of the changes that you calculated.

OpenStudy (anonymous):

A bit late but thanks for the help. xD